# Advances in Radiance Field for Dynamic Scene: From Neural Field to Gaussian Field

### v1

@@ 3.2.4 Representing via Point Tracking
Recovering motion ultimately reduces to establishing reliable point correspondences across time. Early solutions include sparse feature matching and optical flow. Sparse feature matching identifies distinctive keypoints that can be matched across views and has fueled SfM and SLAM pipelines [97]. While effective for camera pose estimation and rigid-scene reconstruction, these sparse correspondences reveal little about dense, non-rigid dynamics. Optical flow extends this by estimating dense 2D correspondences between successive frames [98, 99, 100], but long-range tracking quickly deteriorates under appearance changes, occlusions, or large viewpoint shifts.
To extend temporal support, several works chain short-term matches into longer 2D point trajectories [101, 102, 103]. However, purely image-plane tracking struggles with out-of-plane motions that are more naturally handled in 3D. Scene-flow methods (Sec. 3.2.3) lift correspondences to 3D but restricts motion to adjacent frames.
Recent progress in differentiable volume rendering has enabled dense, long-term 3D tracking driven only by 2D supervision [23, 18, 17]. These approaches jointly optimize a volumetric scene representation and a continuous trajectory field:
@@
where $\mathcal{J}_{\theta}$ returns the 3D position of a point at any time $t$. Trajectories can be realized by composing deformation fields through a canonical space (Eq. 15) or by integrating consecutive flow fields (Eq. 20). Rather than stitching these local matches, several methods fit $\mathcal{J}_{\theta}$ directly over the full sequence, enforcing temporal smoothness and consistency [46, 45]. At the meantime, foundation models such as DINO [104] can also supply robust, semantically rich features that could further stabilize long-range correspondence under severe appearance change. Trajectory-based tracking thus provides a unified motion description in which every scene point follows a continuous 3D path.
@@ 3.2.5 Representing via Factorization
Decomposing high-dimensional, complex signals into lower-dimensional, simpler components represents a fundamental strategy for efficient scene representation. This approach significantly reduces computational complexity while preserving essential scene characteristics. For static neural radiance fields, methods such as TensoRF [5] and EG3D [105] have demonstrated the effectiveness of this principle by representing 3D space using 2D tensors and triplanes, achieving both efficient training and rendering.
When extending to dynamic scenes, the 4D spacetime domain introduces additional complexity that can be effectively managed through factorization techniques. Methods like Hexplane [15] and K-planes [14] decompose this 4D domain into separable components: static spatial features captured in planes $({\mathbf{f}_{xy}},{\mathbf{f}_{yz}},{\mathbf{f}_{xz}})$, and motion-related temporal components represented in planes $({\mathbf{f}_{xt}},{\mathbf{f}_{yt}},{\mathbf{f}_{zt}})$. For any 3D point $\mathbf{x}$ at a specific time $t$, its features are interpolated from these orthogonal planes and processed by a neural network to predict the radiance field properties:
$\mathcal{F}_{\theta}(\mathbf{f}_{xy,yz,zx}(\mathbf{x}),\mathbf{f}_{xt,yt,zt}(\mathbf{x}))\mapsto(\mathbf{c},\sigma).$ (22)
This hyperplane-based factorization effectively disentangles spatial and temporal components, enabling more efficient optimization and better generalization across the temporal domain.
For scenarios where motion is modeled separately—such as in deformation fields or flow fields—a basis-driven decomposition provides an elegant solution. This approach leverages a limited set of shared motion bases $\{\boldsymbol{b}_{i}\}_{i=1}^{B}$ and time-dependent coefficients $c_{i}(t)$ to represent each point’s motion trajectory:
$\mathbf{x}(t)=\sum_{i=1}^{B}c_{i}(t)\,\boldsymbol{b}_{i},\quad\text{with}\quad\boldsymbol{b}_{i}\in\mathbb{R}^{3},\quad c_{i}(t)\in\mathbb{R}$ (23)
@@
The shared basis vectors $\boldsymbol{b}_{i}$ can be implemented as learnable parameters or derived from orthogonal function families such as Fourier or sinusoidal expansions. The coefficients $c_{i}(t)$ dynamically weight each basis vector’s contribution to a point’s overall displacement. Both the basis vectors and coefficients can be jointly optimized with scene geometry and appearance, creating a unified representation that maintains global consistency over time.
This basis-driven approach offers several advantages: it significantly reduces the parameter space required to represent complex dynamic scenes, enforces motion coherence between spatially proximate points, and enables smoother interpolation for novel time steps. By leveraging the underlying structure in real-world motion, these factorization techniques provide a scalable and efficient framework for modeling complex 4D dynamics.
@@ 3.3 Disscussion
Dynamic scene representation presents fundamental challenges, particularly when reconstructing complex motions from limited observations in monocular or sparse view settings. The efficacy of different representation paradigms varies significantly across motion types, with no single approach universally excelling across all scenarios. Rigid motion, efficiently modeled through pose tracking, represents objects using low-dimensional transformations (rotation and translation), making it computationally efficient and interpretable. The key challenge lies in accurately aligning tracking results at different timestamps within a common coordinate system to aggregate temporal information effectively. Articulated motion extends this concept by representing hierarchical relationships between connected rigid parts through kinematic tree structures, balancing strong priors with sufficient flexibility to model entities like humans and animals, though typically requiring category-specific templates that limit broader application.
Non-rigid motion encompasses more general deformations without the structural constraints of rigid or articulated motion, leading to diverse representation strategies with varying strengths. Fig. 5 illustrates the typical representation methods. The 4D spacetime representation directly extends static radiance fields to temporal domain for frame-by-frame optimization, offering conceptual simplicity but lacking explicit point-level motion information and struggling with temporal consistency. Canonical space with deformation fields elegantly disentangles structure from motion by decomposing dynamic scenes into a static canonical space and time-varying deformation fields, though defining appropriate canonical spaces becomes challenging for sequences with large deformations. For longer sequences with significant changes, keyframe-based approaches replace a single canonical space with multiple reference frames, better accommodating topology changes at the cost of increased memory requirements.
Frame-to-frame flow fields model incremental deformations between consecutive frames, naturally accommodating topology changes without requiring global canonical spaces, though they struggle with establishing long-range correspondences and may accumulate errors over time. Point tracking approaches represent each point’s motion as a continuous function across time, establishing dense 3D correspondences but remaining vulnerable to occlusions. Factorization-based representations decompose motion into a limited set of shared basis and time-dependent coefficients, effectively capturing common motion patterns while reducing dimensionality and computational complexity, making complex motion field regularization more intuitive through coefficient regularization.
These representations exist on a spectrum balancing structure and flexibility, with rigid and articulated representations imposing strong priors but sacrificing adaptability, while frame-by-frame approaches maximize flexibility at the cost of temporal consistency. For real-world applications, hybrid approaches combining multiple representation paradigms often yield superior results.
@@ 4.1 Reconstructing with Rigid Motion
Rigid motion reconstruction fundamentally revolves around pose tracking–estimating the 6DoF transformation of each object over time. By tracking the 3D bounding boxes of moving objects, these methods spatially decomposes dynamic scenes into foreground objects and static backgrounds [52], with some methods incorporating dedicated sky modules to handle distant, uncertain regions [106]. Early works such as NSG [25] introduced scene graphs–hierarchical structures where nodes represent individual scene elements (objects or background) and edges encode their spatial relationships as rigid transformation, enabling efficient novel view synthesis of complex dynamic scenes. This concept evolved into multi-level scene graph in ML NSG [110] and dynamic scene graph in ProSGNeRF [109]. Multi-view systems like STaR [32] employ self-supervised tracking alongside scene graphs to jointly optimize object poses and appearances without manual annotations.
Temporal coherence is achieved by linking object instances across frames through consistent pose tracking. Methods like PNF [22] additionally employ meta-learning techniques to initialize category-specific object fields, while HUGS [50] constrains vehicles to ground-plane motion using a unicycle model for improved stability. More recent approaches such as S-NeRF [108] and SG [52] optimize tracked poses jointly with scene parameters, yielding more accurate reconstruction and alignment across frames.
A recent trend is the transition from implicit neural radiance fields (parameterized by MLPs) to explicit 3DGS representations for faster convergence and real-time rendering capabilities. To constrain the ill-posed nature of dynamic scene reconstruction (especially from monocular inputs), methods often leverage auxiliary information such as sementic segmentation [22, 50, 106, 28, 112], optical flow [50, 108], or Lidar [111, 112] and depth data [106] to offer precise boundary, constrain motion space, and resolve scale ambiguity or employ meta-learning techniques to initialize category-specific object fields [22]. UniSim [107] further incorporate feature grids to model finer environmental details while maintaining efficient object-level motion representation.
The practical applications of these approaches span autonomous driving, augmented reality, and robotics. Explicit scene decomposition through rigid motion tracking enables capabilities like object removal [25], trajectory modification [111], and viewpoint manipulation [107]–all crucial for simulation environments and digital twins.
@@ 4.2 Reconstructing with Articulated Motion
Human body: Articulated human body reconstruction has evolved through distinct developmental stages, progressing from simple 2D/3D pose estimation and basic mesh recovery [138] toward photorealistic appearance and fine geometric detail reconstruction. This advancement has been particularly accelerated by the introduction of implicit neural fields [139, 140], especially neural radiance fields [1, 2], which enable high-fidelity modeling of complex human bodies. Early reconstruction approaches strategically integrated parametric human models such as SMPL or SMPL-X as structural priors with neural radiance fields to represent time-varying surface details. These methods established the canonical space paradigm—where the human body is represented in a standardized rest pose while points from the observation space (posed body) are transformed to this canonical reference frame via inverse skinning to query appearance properties like color and density values [117, 115, 116, 35, 141].
@@
The radiance field is supervised using multi-view or monocular images, with the optimization accounting for both appearance and geometry. A fundamental challenge is diffusing skinning weights—originally defined only on the parametric mesh surface—to arbitrary points in 3D space. Solutions range from nearest-neighbor interpolation [114, 142, 83] and barycentric mapping [36] to more sophisticated approaches like Neural Body’s 3D convolutional networks and learnable continuous skinning fields [84, 143, 144]. While early methods defined skinning fields in observation space, these approaches frequently struggle with generalization to novel poses [145, 146]. Recent advances define skinning in canonical space and employ root-finding algorithms to establish bidirectional point correspondences [84, 147, 144, 148, 95], enabling forward skinning transformations that generalize significantly better to out-of-distribution poses and providing more robust reconstruction across complex articulations. Complementary techniques include pixel-aligned features for cross-identity generalization [116, 149, 150, 151], part-wise representations for enhanced detail [148], and hybrid representation for more efficient training and rendering [144, 152, 153].
@@
3D Gaussian Splatting (3DGS) has emerged as a transformative explicit representation for articulated human bodies, offering both quality improvements and dramatic efficiency gains over implicit neural fields [118, 31, 48, 119, 122, 120, 154, 155]. In a typical 3DGS pipeline, Gaussians are initialized based on a parametric template in rest pose and then transformed into observation space via forward skinning, which naturally drives both position and orientation parameters. During this skinning process, the color and density of the Gaussians are typically fixed as initialized to ensure better convergence [49]. This approach fundamentally resolves the correspondence ambiguities present in inverse skinning used by neural implicit representations. Instead of directly optimizing Gaussian parameters, some methods represent Gaussians with learned embeddings, predicting parameters using embeddings sampled via UV mapping [120, 121, 122]. As 3D Gaussian fields lack inherent structure, advanced methods bind Gaussians to structured meshes or tetrahedral cages to enhance animation control and spatial coherence [156, 157]. Several approaches also jointly optimize pose, skinning weights, and skeletal structures alongside Gaussian parameters [158].
@@
Beyond canonical space methods, alternative approaches represent articulated bodies directly in observation space or pursue template-free reconstruction. Direct observation space techniques leverage specialized parameterizations such as UV-based coordinates [159, 36] or skeleton-based local coordinates [113, 114, 160], circumventing the need for explicit canonicalization. Meanwhile, template-free approaches learn articulation parameters—including skeleton structures and skinning weights—entirely from scratch, demonstrating remarkable generalizability across different body types and motion patterns [161, 162, 123]. These methods typically incorporate auxiliary supervision from silhouettes, semantic segmentation, optical flow, or leverage data-driven priors from foundation models like DINO features [104] or CSE embeddings [163] to establish robust correspondences across frames. The integration of 2D computer vision constraints with 3D reconstruction techniques enables increasingly accurate results even from monocular inputs, with recent advances achieving reconstruction times of minutes rather than hours or days, while maintaining high visual quality and animation capabilities.
@@
Hands: Human hands are another typical articulated structure that plays a crucial role in everyday life. Similar to the SMPL body model, MANO [34] represents hand geometry with a pre-defined skeleton, blend shapes, and models hand motion using pose parameters and skinning weights. While MANO provides a widely-used parametric foundation, its relatively coarse mesh has led to developments like MANO-HD [125], which offers high-resolution hand geometry while maintaining compatibility with existing MANO-annotated datasets, enabling personalized hand shape fitting using only monocular RGB videos. For realistic hand appearance, researchers have explored diverse representation strategies including neural fields [124, 164], radiance fields [165, 126], and texture maps [166, 167, 168], with methods like RelightableHands [169], HandRT [170], and URHand [168] enabling physically-based relighting through explicit material modeling. Hand reconstruction approaches increasingly address complex interaction scenarios, including hand-to-hand [165, 171] and hand-to-object interactions [172, 173]. Recent advances have introduced novel representations such as LiveHand’s [126] UVH space parameterization that represents hands relative to the MANO surface without explicit skinning, Nimble’s [174] modeling of inner bones and muscles for enhanced biomechanical realism. MANUS [128] effectively represents articulated hands by utilizing 3D Gaussian Splatting in canonical space and employing forward skinning to convert these Gaussians into posed space, where the model is supervised through multi-view sequences to achieve precise shape and appearance reconstruction. GaussianHand [127] enhances articulated hand modeling by leveraging canonical features to refine blend shapes derived from parametric models like MANO and implementing neural residual skeletons to capture subtle pose-dependent deformations, resulting in more accurate representation of hand poses than methods using only standard linear blend skinning.
Animal: Reconstructing articulated animals presents unique challenges due to the vast diversity of species, making it difficult to adapt a single template to accommodate all morphologies. The SMAL model [175], a pioneering parametric 3D representation, focuses primarily on quadruped animals and provides a foundation for subsequent research. Based on this parametric framework, researchers have demonstrated that animal shapes can be accurately fitted using only 2D image inputs [176]. More recently, neural radiance field approaches have enabled learning detailed geometry and appearance [132, 129, 86], allowing for photorealistic novel view synthesis through volume rendering. These methods typically employ a dual-level representation strategy: a category-level template for general morphology combined with instance-level corrections to capture individual variations [87].
Several innovative approaches have moved beyond template meshes, instead using only skeletal structures as priors [129, 130, 177], where the posed animal is driven entirely by skeletal motion. Even more remarkably, some methods operate without any species-specific prior knowledge, learning animal models automatically from raw data [86, 131, 178], demonstrating the versatility of articulated motion representations across diverse morphologies. For example, BANMo [86] learns neural bones and skinning weights directly from casual videos, representing animals as density and color radiance fields in canonical space, with point correspondences established through bidirectional neural skinning operations. ARTEMIS [129] represents geometry and appearance using neural radiance feature voxel grids, with posed animals warped via skeletal motion and rendered through neural rendering techniques. CoP3D [131] utilizes pixel-aligned features to estimate density and color without explicitly modeling motion. Learning the articulation from casually monocular videos remains an inherently ill-posed problem, these methods leverage various auxiliary information sources to constrain the solution space, including foreground masks [132, 129], optical flow [86, 131], surface normals [87], and semantically rich features from foundation models such as CSE [132] and DINO [177].
Objects: Unlike human or animal bodies with consistent skeletal structures, general articulated objects (such as laptops, scissors, and other mechanical devices) present unique reconstruction challenges due to their diverse topologies and joint configurations. For these objects, using pre-defined kinematic trees becomes impractical, as each object type features distinct articulation patterns. The key challenge in reconstructing such general articulated objects involves three interrelated tasks: accurately segmenting the constituent parts, defining appropriate joint motion types (e.g., rotational, prismatic), and estimating precise joint motion parameters for each articulated state [179].
@@
NeRFlow [38] ICCV’21 monocular flow field MLP
@@
DCT-NeRF [23] Arxiv’21 monocular TV/ARAP/cycle/$L_{1}$ trajectory MLP
PVG [46] Arxiv’23 monocular $L_{1}$ trajectory 3DGS

### v2

@@ 3.2.4 Representing via Point Tracking
Recovering motion ultimately reduces to establishing reliable point correspondences across time. Early solutions include sparse feature matching and optical flow. Sparse feature matching identifies distinctive keypoints that can be matched across views and has fueled SfM and SLAM pipelines [86]. While effective for visual localization, these sparse correspondences reveal little about dense, non-rigid dynamics. Optical flow extends this by estimating dense 2D correspondences between successive frames [87, 88, 89], but long-range tracking quickly deteriorates under appearance changes, occlusions, or large viewpoint shifts. several works chain short-term matches into longer 2D point trajectories [90, 91, 92]. However, purely image-plane tracking struggles with out-of-plane motions that are more naturally handled in 3D.
Recent progress in differentiable rendering and radiance field has enabled dense, long-term 3D tracking driven by only 2D supervision [23, 18, 17]. These approaches jointly optimize a scene volume and a continuous trajectory field:
@@
where $\mathcal{J}_{\theta}$ describes 3D trajectory of any point in 3D space. Rather than stitching local matches via Eq. 13 or Eq. 17, fitting such per-point trajectory $\mathcal{J}_{\theta}$ directly over the full sequence as a whole enforces temporal consistency [46, 45].
@@ 3.2.5 Representing via Factorization
Decomposing high-dimensional, complex signals into lower-dimensional, simpler components represents a fundamental strategy for signal processing, which is also applicable in motion analysis. This approach significantly reduces computational complexity while preserving essential motion characteristics. For static neural radiance fields, methods such as TensoRF [5] and EG3D [93] have demonstrated the effectiveness of this principle by representing 3D volume via 2D tensors or triplanes, achieving both efficient training and rendering.
When extending to dynamic scenes, the 4D spacetime introduces additional complexity that can be effectively managed through factorization techniques. Typically, this 4D domain could be decomposed into separable components: static spatial features captured in $({\mathbf{f}_{xy}},{\mathbf{f}_{yz}},{\mathbf{f}_{xz}})$ planes, and motion-related temporal components represented in $({\mathbf{f}_{xt}},{\mathbf{f}_{yt}},{\mathbf{f}_{zt}})$ planes. For any 3D point $\mathbf{x}$ at a specific time $t$, its features are interpolated from these orthogonal hyperplanes and processed by a neural network to predict the radiance field properties:
$\mathcal{F}_{\theta}(\mathbf{f}_{xy,yz,zx}(\mathbf{x}),\mathbf{f}_{xt,yt,zt}(\mathbf{x}))\longmapsto(\mathbf{c},\sigma).$ (19)
This hyperplane-based factorization effectively disentangles spatial and temporal components, enabling more efficient optimization and better generalization.
For scenarios where motion is modeled separately, such as in deformation fields or flow fields, a basis-driven decomposition provides an elegant solution. This approach leverages a limited set of shared motion bases $\{\mathbf{b}_{i}\}_{i=1}^{B}$ and time-dependent coefficients $c_{i}(t)$ to represent each point’s motion trajectory:
$\mathbf{x}(t)=\sum_{i=1}^{B}c_{i}(t)\,\mathbf{b}_{i},\quad\text{with}\ \mathbf{b}_{i}\in\mathbb{R}^{3},\ c_{i}(t)\in\mathbb{R}$ (20)
@@
The shared basis vectors $\mathbf{b}_{i}$ can be implemented as learnable parameters or derived from orthogonal function families such as Fourier or sinusoidal expansions. The coefficients $c_{i}(t)$ dynamically weight each basis vector’s contribution to a point’s overall movement. Both the basis vectors and coefficients can be jointly optimized with scene geometry and appearance, creating a compact motion representation that maintains global consistency over time.
Figure 4: Illustration of typical motion representation methods.
@@ 3.3 Disscussion
Dynamic motion representation methods span a spectrum balancing structural constraints and flexibility, with each approach offering distinct trade-offs, illustrated in Fig. 4. Rigid motion representations efficiently model objects through 6DoF pose tracking (sometimes incorporating scale factors for distance variations), providing computational efficiency and interpretability while inherently limiting complex deformations. Articulated motion extends this concept by modeling hierarchical relationships between connected rigid parts through kinematic chains, effectively representing entities like human bodies while typically requiring category-specific templates as prior knowledge. Canonical space with deformation fields elegantly disentangles geometry and appearance from motion by decomposing scenes into a static canonical space and time-varying deformation fields, enabling sophisticated motion analysis and canonical manipulation, though defining appropriate canonical spaces becomes challenging for sequences with large deformations or extended temporal spans.
4D spacetime representations directly extend static scene modeling to the temporal domain for frame-by-frame optimization, offering conceptual simplicity but lacking explicit point correspondence and struggling with cross-view consistency. Frame-to-frame flow fields model incremental deformations between consecutive frames, naturally accommodating topology changes without requiring global canonical spaces, though they accumulate errors over extended sequences. Point tracking approaches represent trajectories as continuous functions across time, establishing dense 3D correspondences while remaining vulnerable to occlusions. Factorization-based representations decompose motion into shared basis functions and time-dependent coefficients, significantly reducing parameter dimensionality while enforcing motion coherence between spatially proximate points. This representation spectrum reflects fundamental trade-offs: rigid and articulated approaches impose strong priors but sacrifice adaptability, while frame-by-frame methods maximize flexibility at the cost of temporal consistency, suggesting that hybrid approaches combining complementary paradigms often yield superior results for complex real-world applications.
@@ 4.1 Reconstructing with Rigid Motion
Rigid motion reconstruction fundamentally revolves around pose tracking, estimating the 6DoF transformation of each object over time. By tracking the 3D bounding boxes of moving objects, these methods spatially decomposes dynamic scenes into foreground objects and static backgrounds [52], with some methods incorporating dedicated sky modules to handle distant, uncertain regions [94]. Early works such as NSG [25] introduced the scene graph, a hierarchical structure where nodes represent individual scene elements (objects or background) and edges encode their spatial relationships as rigid transformation, enabling efficient novel view synthesis of complex dynamic scenes. This concept evolved into multi-level scene graph in ML NSG [97] and dynamic scene graph in ProSGNeRF [100]. Multi-view systems like STaR [32] employ self-supervised tracking alongside scene graph to jointly optimize the object poses without manual annotations.
Temporal coherence is achieved by linking object instances across frames through consistent pose tracking. Methods like PNF [22] additionally employ meta-learning techniques to initialize category-specific object fields, while HUGS [50] constrains vehicles to ground-plane motion using a unicycle model for improved stability. More recent approaches, such as S-NeRF [96] and SG [52], optimize tracked poses jointly with scene parameters, yielding more accurate reconstruction and alignment across frames.
A recent trend is the transition from implicit neural radiance fields (parameterized by MLPs) to explicit 3DGS representations for faster convergence and real-time rendering capabilities. To constrain the ill-posed nature of dynamic scene reconstruction (especially from monocular inputs), methods often leverage auxiliary information such as sementic segmentation [22, 50, 94, 28, 99], optical flow [50, 96], or Lidar [98, 99] and depth data [94] to offer precise boundary, constrain motion space, and resolve scale ambiguity. UniSim [95] further incorporate feature grids to model finer environmental details while maintaining efficient object-level motion representation. The practical applications of these approaches span autonomous driving, augmented reality, and robotics. Explicit scene decomposition through rigid motion tracking enables capabilities like object removal [25], trajectory modification [98], and viewpoint manipulation [95], all crucial for simulation environments and digital twins.
@@ 4.2 Reconstructing with Articulated Motion
Human Body. Articulated human body reconstruction has evolved through distinct developmental stages, progressing from simple 2D/3D pose estimation and basic mesh recovery [128] toward photorealistic appearance and fine geometric detail reconstruction. This advancement has been particularly accelerated by the introduction of implicit neural fields [129, 130], especially recent radiance fields [1, 2]. Early reconstruction approaches strategically integrated parametric human models such as SMPL or SMPL-X as structural priors with neural radiance fields to represent time-varying surface details. These methods established the canonical space paradigm, where the human body is represented in a rest pose while points from the observation space (posed body) are transformed to this canonical reference via inverse skinning to query field properties like color and density values [105, 103, 104, 35, 131].
@@
The learning process is supervised using multi-view or monocular images, with the optimization accounting for both appearance and geometry. A fundamental challenge is diffusing skinning weights to arbitrary points in 3D space. Solutions range from nearest-neighbor interpolation [102, 132, 75] and barycentric mapping [36] to more sophisticated approaches like Neural Body’s 3D convolutional networks and learnable continuous skinning fields [76, 133, 134]. While early methods defined skinning fields in observation space, these approaches frequently struggle with generalization to novel poses [135, 136]. Recent advances define the skinning field in canonical space and employ root-finding algorithms to establish bidirectional point correspondences [137, 134, 138, 106], enabling forward skinning transformations that generalize significantly better to out-of-distribution poses and providing more robust reconstruction across complex articulations. Complementary techniques include pixel-aligned features for cross-identity generalization [104, 139, 140, 141], part-wise representations for enhanced detail [138], and hybrid representation for more efficient training and rendering [134, 142, 143].
@@
3DGS has emerged as a transformative explicit representation for articulated human bodies, offering both quality improvements and dramatic efficiency gains over implicit neural fields [107, 31, 48, 108, 111, 109, 144, 145]. In a typical 3DGS pipeline, Gaussians are initialized based on a parametric template in rest pose and then transformed into observation space via forward skinning, which naturally drives both position and orientation parameters. During this skinning process, the color and density of the Gaussians are typically fixed as initialized to ensure better convergence [49]. This scheme fundamentally resolves the correspondence ambiguities present in inverse skinning used by neural implicit representations. Instead of directly optimizing Gaussian parameters, some methods represent Gaussians with learned embeddings, predicting parameters using embeddings sampled via UV mapping [109, 110, 111]. As 3D Gaussian fields lack inherent structure, advanced methods bind Gaussians to structured meshes or tetrahedral cages to enhance animation control and spatial coherence [146, 147]. Several approaches also jointly optimize pose, skinning weights, and skeleton alongside Gaussian parameters [148].
@@
Beyond canonical space methods, alternative approaches represent articulated bodies directly in observation space or pursue template-free reconstruction. Direct observation space techniques leverage specialized parameterizations such as UV-based coordinates [149, 36] or skeleton-based local coordinates [101, 102, 150], circumventing the need for explicit canonicalization. Meanwhile, template-free approaches learn articulation parameters, including skeleton structures and skinning weights, entirely from scratch, demonstrating remarkable generalizability across different body types and motion patterns [151, 152, 112]. These methods typically incorporate auxiliary supervision from silhouettes, semantic segmentation, optical flow, or leverage data-driven priors from foundation models like DINO features [153] or CSE embeddings [154] to establish robust correspondences across frames.
@@
Hands. Human hands are another typical articulated structure that plays a crucial role in everyday life. Similar to the SMPL body model, MANO [34] represents hand geometry with a pre-defined skeleton, blend shapes, and models hand motion using pose parameters and skinning weights. While MANO provides a widely-used parametric foundation, its relatively coarse mesh has led to developments like MANO-HD [114], which offers high-resolution hand geometry while maintaining compatibility with existing MANO-annotated datasets. For realistic hand appearance, researchers have explored diverse representation strategies including neural fields [113, 155], radiance fields [156, 115], and texture maps [157, 158, 159], with methods like RelightableHands [160], HandRT [161], and URHand [159] enabling physically-based relighting through explicit material modeling. Hand reconstruction approaches increasingly address complex interaction scenarios, including hand-to-hand [156, 162] and hand-to-object interactions [163, 164]. Recent advances have introduced novel representations such as LiveHand’s [115] UVH space parameterization that represents hands relative to the MANO surface without explicit skinning, Nimble’s [165] modeling of inner bones and muscles for enhanced biomechanical realism. MANUS [117] effectively represents articulated hands by utilizing 3D Gaussian Splatting in canonical space and employing forward skinning to convert these Gaussians into posed space, where the model is supervised through multi-view sequences to achieve precise shape and appearance reconstruction. GaussianHand [116] enhances articulated hand modeling by leveraging canonical features to refine blend shapes derived from parametric models like MANO and implementing neural residual skeletons to capture subtle pose-dependent deformations, resulting in a more accurate representation of hand poses than methods using only standard linear blend skinning.
@@
NeRFlow [38] ICCV’21 monocular scene flow MLP
@@
Animal. Reconstructing articulated animals presents unique challenges due to the vast diversity of species, making it difficult to adapt a single template to accommodate all morphologies. The SMAL model [176], a pioneering parametric model primarily for quadruped animals, provides a foundation for subsequent research. Based on this parametric framework, researchers have demonstrated that animal shapes can be accurately fitted using only 2D image inputs [177]. More recently, neural radiance field approaches have enabled learning detailed geometry and appearance [122, 118, 119], allowing for photorealistic novel view synthesis through volume rendering. These methods typically employ a dual-level representation: a category-level template for general morphology combined with instance-level corrections to capture individual time-varying variations [178].
Several innovative approaches have moved beyond template meshes, instead using only skeletal structures as priors [118, 120, 179], where the posed animal is driven entirely by skeletal motion. Even more remarkably, some methods operate without any species-specific prior knowledge, learning animal models automatically from raw data [119, 121, 180], demonstrating the versatility of articulated motion representations across diverse morphologies. For example, BANMo [119] learns neural bones and skinning weights directly from casual videos, representing animals as a neural radiance field in canonical space, with point correspondences established through bidirectional neural skinning operations. ARTEMIS [118] represents geometry and appearance using neural feature voxel grids, with posed animals warped via skeletal motion and rendered through neural rendering techniques. CoP3D [121] utilizes pixel-aligned features to estimate density and color without explicitly modeling motion. Learning the articulation from casually monocular videos remains an inherently ill-posed problem, these methods also leverage various auxiliary information sources to constrain the solution space, including foreground masks [122, 118], optical flow [119, 121], surface normals [178], and semantically rich features from foundation models such as CSE [122] and DINO [179].
Objects. Unlike human or animal bodies with consistent skeletal structures, general articulated objects (such as laptops, scissors, and other mechanical devices) present unique reconstruction challenges due to their diverse topologies and joint configurations. For these objects, using pre-defined kinematic trees becomes impractical, as each object type features distinct articulation patterns. The key challenge in reconstructing such general articulated objects involves three interrelated tasks: accurately segmenting the constituent parts, defining appropriate joint motion types (e.g., rotational, prismatic), and estimating precise joint motion parameters for each articulated state [181].